OpenStudy (anonymous):

calculate the time to pass a 200 m platform for a 200m long train having 90kmh-1 strady speed .. solve>

OpenStudy (chmvijay):

@John_ES can u help her:)

OpenStudy (yttrium):

Just be reminded that \[t = d/s\]

OpenStudy (yttrium):

Do you know what to do now @amelykatlyn ?

OpenStudy (chmvijay):

@mathslover can u help her

OpenStudy (anonymous):

Total distance covered by train will be (length of train + length of platform) total distance = 400m speed 90 km/h

OpenStudy (anonymous):

iether you change 400 m in km and find time or the second way is to find speed in m/s then find time

OpenStudy (anonymous):

answer will be 16 sec

OpenStudy (anonymous):

change km/h in m/sec \[(90\times 1000 m)\div3600 \sec = 25 m/s\] time =distance/speed 400/25= 16 sec